Question 13

Cullumber Company purchased equipment on March 27, 2018, at a cost of $328,000. Management is contemplating the merits of using the diminishing-balance or units-of-production method of depreciation instead of the straight-line method, which it currently uses for other equipment. The new equipment has an estimated residual value of $8,000 and an estimated useful life of either four years or 80,000 units. Demand for the products produced by the equipment is sporadic so the equipment will be used more in some years than in others. Assume the equipment produces the following number of units each year: 14,600 units in 2018; 20,400 units in 2019; 20,000 units in 2020; 20,000 units in 2021; and 5,000 units in 2022. Cullumber has a December year end.

Collapse question part

(a)

Prepare separate depreciation schedules for the life of the equipment using: (Round depreciation per unit to 2 decimal places, e.g. 5.28 and final answers to 0 decimal places, e.g. 5,275.)

Straight-line method:
Year Depreciable
Amount
Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
$

2018 $

$

$

2019

2020

2021

2022


Double-diminishing-balance method:
Year Opening
Carrying
Amount
Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
$

2018 $

$

$

2019

2020

2021

2022


Units-of-production method:
Year Units-of-Production Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
$

2018

$

$

2019

2020

2021

2022
Answer #1
1) Straight line depreciation = (cost -salvage value)/life of assets
depreciation = (328000-8000)/4 = 80000
depreciation exp accumulated depreciation book value
0 328000
2018 80000*9/12 60000 60000 268000
2019 80000 140000 188000
2020 80000 220000 108000
2021 80000 300000 28000
2022 80000*3/12 20000 320000 8000
2) Double declining method
depreciation rate = (328000/4)/328000*2 = 50%
depreciation exp accumulated depreciation book value
328000
2018 328000*50%*9/12 123000 123000 205000
2019 205000*50% 102500 225500 102500
2020 102500*50% 51250 276750 51250
2021 51250*50% 25625 302375 25625
2022 25625*50%*3/12 3203 305578 22422
3) Unit of production method
depreciation = (cost-salvage value)/total machine hours
depreciation per machine hr = (328000-8000)/80000 = 4
depreciation exp accumulated depreciation book value
2018 328000
2018 14600*4 58400 58400 269600
2019 20400*4 81600 140000 188000
2020 20000*4 80000 220000 108000
2021 20000*4 80000 300000 28000
2022 5000*4 20000 320000 8000